OpenStudy (anonymous):

Algebra 2 please help

OpenStudy (anonymous):

find the logarithm

OpenStudy (anonymous):

of?

OpenStudy (anonymous):

\[\log_{1/4} (1/64)\]

OpenStudy (anonymous):

your job is to solve \[\left(\frac{1}{4}\right)^x=\frac{1}{64}\] for \(x\)

OpenStudy (anonymous):

since \(4^3=64\) you have \[\left(\frac{1}{4}\right)^3=\frac{1}{64}\] and therefore \[\log_{\frac{1}{4}}(\frac{1}{64})=3\]

OpenStudy (anonymous):

thank you for your help :)

OpenStudy (anonymous):

yw hope it was more or less clear

OpenStudy (anonymous):

it was really clear thank you